OpenStudy (anonymous):

Consider the functions f(x) = 3x^2 + 1 and g(x) = x + 3. What is the value of f[g(−1)]? A. 7 B. 37 C. 13 D. 12

OpenStudy (anonymous):

13, c

OpenStudy (anonymous):

wat

OpenStudy (lgbasallote):

heh i can say D..doesnt make me any more wrong :)

OpenStudy (anonymous):

wait i need to fix something

OpenStudy (anonymous):

there

OpenStudy (lgbasallote):

first solve for g(-1). i assume you know how to?

OpenStudy (anonymous):

could you wlk me through it

OpenStudy (lgbasallote):

to find g(-1) simply replace the x of g(x) with -1 g(x) = x + 3 g(-1) = -1 + 3

OpenStudy (lgbasallote):

so what's g(-1)

OpenStudy (anonymous):

2

OpenStudy (lgbasallote):

right. now since g(-1) = 2 f(g(-1) ) = f(2) agree?

OpenStudy (anonymous):

agree

OpenStudy (lgbasallote):

so now solve for f(2)

OpenStudy (anonymous):

thinking

OpenStudy (anonymous):

the answer is 13 after all

OpenStudy (anonymous):

lol the person got the answer so quick

OpenStudy (lgbasallote):

lucky guess lol